# Faster Linear Systems and Matrix Norm Approximation via Multi-level Sketched Preconditioning

### v1

@@ 6 Analysis of Inexact Preconditioned Lanczos Iteration
We address this issue by proving a new result for the Lanczos method with inexact preconditioning that maintains the accelerated rate. We suspect that a similar result can be achieved for PCG, which is more commonly used in practice than Lanczos, although we leave this effort to future work. We also note that it might be possible to apply other approaches to obtain accelerated rates. For example, [ST14] give an analysis of inexact preconditioned Chebyshev iteration. However, their analysis requires that the approximation algorithm used to apply $\mathbf{M}^{-1}$ is a linear operator, and Chebyshev iteration is both more difficult to use than Lanczos, requiring eigenvalue bounds, and converges more slowly in practice (even though it has the same asymptotic guarantees) [ACG+23].
@@ 6.1 Inexact Preconditioned Lanczos Method
16:return $\mathbf{x}_{t}^{\prime}=z^{\prime}\overline{\mathbf{Q}}\mathbf{T}^{-1}\mathbf{e}_{1}$ where $\mathbf{e}_{1}\in\mathbb{R}^{t}$ denotes the first standard basis vector.
@@
$\|\mathbf{x}_{t}-\mathbf{A}^{-1}\mathbf{b}\|_{\mathbf{A}}$ $\leq\frac{\epsilon{\lambda_{1}}}{\kappa_{\mathbf{M}}\lambda_{n}}\|\mathbf{A}^{-1}\mathbf{b}\|_{\mathbf{A}}=\epsilon\|\mathbf{A}^{-1}\mathbf{b}\|_{\mathbf{A}}.\qed$
@@ 7 Lower Bound for Linear Systems with $k$ Large Singular Values
$\mathcal{A}=\Big{\{}\mathbf{I}_{n}+\mathbf{e}_{i}^{\top}\mathbf{e}_{j}-\mathbf{e}_{j}\mathbf{e}_{i}^{\top}\quad:\quad i\neq j\Big{\}}.$
@@
$\mathbf{A}^{\top}\mathbf{A}=(\mathbf{I}_{n}+\mathbf{e}_{i}^{\top}\mathbf{e}_{j}-\mathbf{e}_{j}\mathbf{e}_{i}^{\top})^{\top}(\mathbf{I}_{n}+\mathbf{e}_{i}^{\top}\mathbf{e}_{j}-\mathbf{e}_{j}\mathbf{e}_{i}^{\top})=\mathbf{I}_{n}+\mathbf{e}_{i}\mathbf{e}_{i}^{\top}+\mathbf{e}_{j}\mathbf{e}_{j}^{\top}.$

### v2

@@ 6 Analysis of Inexact Preconditioned Lanczos Iteration
We address this issue by proving a new result for the Lanczos method with inexact preconditioning that maintains the accelerated rate. We suspect that a similar result can be achieved for PCG, which is more commonly used in practice than Lanczos, although we leave this effort to future work. As discussed in Section 1.2, we also note that our analysis of Lanczos is not strictly necessary for this paper. For example, [ST14b] give an analysis of inexact preconditioned Chebyshev iteration, which could have been used instead. However, the preconditioned Lanczos method is easier to implement (e.g., does not require eigenvalue bounds) and much more widely used in practice since it tends to converge much more quickly than Chebyshev iteration, even though it has the same worst-case guarantees [ACG+23]. As such, we expect that our stability analysis of the preconditioned Lanczos method will be of general interest beyond this work.
@@ 6.1 Inexact Preconditioned Lanczos Method
16:return $\mathbf{x}_{t}^{\prime}=z^{\prime}\overline{\mathbf{Q}}\mathbf{T}^{\prime-1}\mathbf{e}_{1}$ where $\mathbf{e}_{1}\in\mathbb{R}^{t}$ denotes the first standard basis vector.
@@
$\|\mathbf{x}_{t}-\mathbf{A}^{-1}\mathbf{b}\|_{\mathbf{A}}$ $\leq\frac{\epsilon{\lambda_{1}}}{\kappa_{\mathbf{M}}\lambda_{n}}\|\mathbf{A}^{-1}\mathbf{b}\|_{\mathbf{A}}=\epsilon\|\mathbf{A}^{-1}\mathbf{b}\|_{\mathbf{A}}.$
@@ 7 Lower Bound for Linear Systems with $k$ Large Singular Values
$\mathcal{A}=\Big{\{}\mathbf{I}_{n}+\mathbf{e}_{i}\mathbf{e}_{j}^{\top}-\mathbf{e}_{j}\mathbf{e}_{i}^{\top}\quad:\quad i\neq j\Big{\}}.$
@@
$\mathbf{A}^{\top}\mathbf{A}=(\mathbf{I}_{n}+\mathbf{e}_{i}\mathbf{e}_{j}^{\top}-\mathbf{e}_{j}\mathbf{e}_{i}^{\top})^{\top}(\mathbf{I}_{n}+\mathbf{e}_{i}\mathbf{e}_{j}^{\top}-\mathbf{e}_{j}\mathbf{e}_{i}^{\top})=\mathbf{I}_{n}+\mathbf{e}_{i}\mathbf{e}_{i}^{\top}+\mathbf{e}_{j}\mathbf{e}_{j}^{\top}.$
@@
Acknowledgements
MD would like to acknowledge NSF CAREER for partial support. CM was partially supported by NSF Award (2045590).